# A Simulation Based Inference Approach to the Dynamics of the MW–LMC System - Validation

### v1

@@ 6.2 Choice of prior distributions
To assess whether our inference of the MW and LMC masses are being biased by the choice of informative prior, we run a smaller set of $4,096$ simulations with the MW and LMC masses drawn from uninformative, flat, prior distributions. These are a distinct sample of simulations from those used throughout the main body of this work and are only used in this section. Specifically we sample the masses from $M_{200,\rm MW}\sim\,\mathcal{U}(1,100)\times 10^{11}M_{\odot}$ and $M_{\rm LMC}\sim\,\mathcal{U}(2.5,100)\times 10^{10}M_{\odot}$. We then randomly select one of these rigid simulations from our sample and we aim to recover its true parameters. This test exactly follows the methodology of Sec. 4.1. The only difference being that the MW and LMC masses were drawn from uninformative rather than informative prior distributions. As before, we estimate the MW and LMC mass posterior distributions with knowledge of the LMC present day position and velocity and MW reflex motion parameters. We find that the true simulation mass values fall well within the $16^{\rm th}-84^{\rm th}$ percentiles of the posterior. This indicates that, in the case where we supply reflex motion information as data points, our SBI framework is not subject to any confirmation biasing of the MW and LMC mass by using informative prior distributions.
@@ 6.3 Cosmological MW–LMC analogues
Our validation test to use the presented SBI formalism to capture the complexity of a cosmological simulation analogue of the MW–LMC system in Sec. 4.3 focused only on a single analogue; FIRE Latte m12b. Indeed, many other MW–LMC analogues have been identified within cosmological simulations e.g., there are more within the FIRE Latte suite itself (Wetzel et al., 2023), plus many more in the Milky Way-est (Buch et al., 2024), the DREAMS (Rose et al., 2025) and the Auriga simulation suites (Grand et al., 2017, 2024). As the rotation to the real MW–LMC orientation (Arora et al., 2024, appendix A) can be easily generalised, this implies that we could apply our SBI framework to any cosmological MW–LMC analogue. This may be beneficial to start to build up statistics on e.g, the inferred reflex motion parameters of MW analogues for a variety of unique LMC analogue masses and infall scenarios. Or, this could even allow the existing sets of cosmological MW–LMC analogues to be used as the input simulations for the application of SBI to observational data.
@@ 6.4 Application to observational data
For the dynamical quantities considered in this work, there already exists observational measurements. For example, there is the DESI BHB (Byström et al., 2024) and also the all-sky red giant stars out to $160\,\rm kpc$ from the combined H3 + SEGUE + MagE dataset (Chandra et al., 2025). Indeed, they measure the MW reflex motion parameters $v_{\rm travel},\,l_{\rm apex}$ and $b_{\rm apex}$, but they also report on the measurement of the Galactocentric solar motion corrected stellar radial velocities, $v_{\rm GSR}$. In both sets of observations, stars in the southern Galactic hemisphere exhibit negative mean radial velocities, i.e., $\langle v_{\rm GSR}\rangle<0\,\rm km\,\rm s^{-1}$, due to the perturbation from the LMC. Interestingly, the same measurement for stars in the northern hemisphere disagrees between the studies. The DESI sample in Byström et al. (2024) shows that radial velocities remain consistent with the equilibrium expectation of $\langle v_{\rm GSR}\rangle\sim 0\,\rm km\,\rm s^{-1}$, whereas the H3 + SEGUE + MagE sample in (Chandra et al., 2025) displays a positive mean radial velocity but with the opposite sign i.e., $\langle v_{\rm GSR}\rangle>0\,\rm km\,\rm s^{-1}$. This is an intriguing discrepancy between the two surveys, and it is one which we aim to address with the application of our SBI framework. In our Figs. 1 & 3 & 4 we show the on-sky $\langle v_{\rm GSR}\rangle$ distributions for the corresponding MW–LMC simulation stellar halo using stars beyond $50\,\rm kpc$ and subtracting off the mean motion across the entire sky. Generally, the strength of the average radial velocity signal will depend on the model for the MW and LMC. Plus, the boundaries of any survey’s sky coverage could influence the results of averaging in each hemisphere e.g., the northern Galactic hemisphere in all of Figs. 1 & 3 & 4 contains regions of both positive and negative average radial velocity. The existing discrepancies for the observational results for the reflex motion and the solar motion corrected stellar radial velocities (Byström et al., 2024; Yaaqib et al., 2024; Chandra et al., 2025) motivate us to apply our SBI framework to this problem in order to better understand these dynamical quantities.
@@ 7 Conclusions
    Our SBI framework can upscale the fidelity of simpler simulations to more complex ones. This allows rapid exploration of large model parameters spaces at the fraction of the computational cost required to run the higher fidelity simulations.
@@
Overall, we have successfully validated using our SBI framework to understand MW dynamics across rigid, deforming, and cosmological MW–LMC models. This motivates us to apply our methodology to real MW stellar halo data from the DESI (Byström et al., 2024) and H3 + SEGUE + MagE datasets (Chandra et al., 2025) to better understand properties of the MW, LMC, and the induced MW reflex motion.
@@ Acknowledgements
Software: sbi (Tejero-Cantero et al., 2020), agama (Vasiliev, 2019), gala (Price-Whelan, 2017), NumPy (Harris et al., 2020), Matplotlib (Hunter, 2007), Seaborn (Waskom, 2021), corner (Foreman-Mackey, 2016), Astropy (Astropy Collaboration et al., 2013, 2018, 2022), SciPy (Virtanen et al., 2020).

### v2

@@ 6.2 Choice of prior distributions
To assess whether our inference of the MW and LMC masses are being biased by the choice of informative prior, we run a distinct set of $4,096$ simulations with the MW and LMC masses drawn from uninformative, flat, prior distributions. These are a distinct sample of simulations from those used throughout the main body of this work and are only used in this section. Specifically we sample the masses from $M_{200,\rm MW}\sim\,\mathcal{U}(1,100)\times 10^{11}M_{\odot}$ and $M_{\rm LMC}\sim\,\mathcal{U}(2.5,100)\times 10^{10}M_{\odot}$. We then randomly select one of these rigid simulations from our sample and we aim to recover its true parameters. This test exactly follows the methodology of Sec. 4.1. The only difference being that the MW and LMC masses were drawn from uninformative rather than informative prior distributions. As before, we estimate the MW and LMC mass posterior distributions with knowledge of the LMC present day position and velocity and MW reflex motion parameters. We find that the true simulation mass values fall well within the $16^{\rm th}-84^{\rm th}$ percentiles of the posterior. This indicates that, in the case where we supply reflex motion information as data points, our SBI framework is not subject to any confirmation biasing of the MW and LMC mass by using informative prior distributions.
@@ 6.3 Cosmological MW–LMC analogues
Our validation test to use the presented SBI formalism to capture the complexity of a cosmological simulation analogue of the MW–LMC system in Sec. 4.3 focused only on a single analogue; FIRE Latte m12b. Indeed, many other MW–LMC analogues have been identified within cosmological simulations e.g., there are more within the FIRE Latte suite itself (Wetzel2023), plus many more in the Milky Way-est (Buch2024), the DREAMS (Rose2025) and the Auriga simulation suites (Grand2017; Grand2024). As the rotation to the real MW–LMC orientation (Arora2024, appendix A) can be easily generalised, this implies that we could apply our SBI framework to any cosmological MW–LMC analogue. This may be beneficial to start to build up statistics on e.g, the inferred reflex motion parameters of MW analogues for a variety of unique LMC analogue masses and infall scenarios. Or, this could even allow the existing sets of cosmological MW–LMC analogues to be used as the input simulations for the application of SBI to observational data.
@@ 6.4 Application to observational data
For the dynamical quantities considered in this work, there already exists observational measurements. For example, there is the DESI BHB (Bystrom2025) and also the all-sky red giant stars out to $160\,\rm kpc$ from the combined H3 + SEGUE + MagE dataset (Chandra2025). Indeed, they measure the MW reflex motion parameters $v_{\rm travel},\,l_{\rm apex}$ and $b_{\rm apex}$, but they also report on the measurement of the Galactocentric solar motion corrected stellar radial velocities, $v_{\rm GSR}$. In both sets of observations, stars in the southern Galactic hemisphere exhibit negative mean radial velocities, i.e., $\langle v_{\rm GSR}\rangle<0\,\rm km\,\rm s^{-1}$, due to the perturbation from the LMC. Interestingly, the same measurement for stars in the northern hemisphere disagrees between the studies. The DESI sample in Bystrom2025 shows that radial velocities remain consistent with the equilibrium expectation of $\langle v_{\rm GSR}\rangle\sim 0\,\rm km\,\rm s^{-1}$, whereas the H3 + SEGUE + MagE sample in (Chandra2025) displays a positive mean radial velocity i.e., $\langle v_{\rm GSR}\rangle>0\,\rm km\,\rm s^{-1}$. This is an intriguing discrepancy between the two surveys, and it is one which we have addressed with the application of our SBI framework to these datasets (see, Brooks2025). In our Figs. 1 & 3 & 4 we show the on-sky $\langle v_{\rm GSR}\rangle$ distributions for the corresponding MW–LMC simulation stellar halo using stars beyond $50\,\rm kpc$ and subtracting off the mean motion across the entire sky. Generally, the strength of the average radial velocity signal will depend on the model for the MW and LMC. Plus, the boundaries of any survey’s sky coverage could influence the results of averaging in each hemisphere e.g., the northern Galactic hemisphere in all of Figs. 1 & 3 & 4 contains regions of both positive and negative average radial velocity. The existing discrepancies for the observational results for the reflex motion and the solar motion corrected stellar radial velocities (Bystrom2025; Yaaqib2024; Chandra2025) motivate us to apply our SBI framework to this problem in order to better understand these dynamical quantities (Brooks2025).
@@ 7 Conclusions
    Our SBI framework trained on many low fidelity MW–LMC simulations retains enough of the relevant physics of the higher fidelity simulations to avoid model misspecification. This is promising for the application to real datasets as it allows for the rapid exploration of large model parameter spaces at a fraction of the computational cost required to run the higher fidelity simulations.
@@
Overall, we have successfully validated our SBI framework to understand the MW dynamics across rigid, deforming, and cosmological MW–LMC models. This motivates us to apply our methodology to real MW stellar halo data from the DESI (Bystrom2025) and H3 + SEGUE + MagE datasets (Chandra2025) to better understand properties of the MW, LMC, and the induced MW reflex motion (Brooks2025).
@@ Acknowledgements
Software: sbi (tejero-cantero2020sbi), agama (2019MNRAS.482.1525V), gala (gala), NumPy (harris2020array), Matplotlib (Hunter:2007), Seaborn (Waskom2021), corner (corner), Astropy (astropy:2013; astropy:2018; astropy:2022), SciPy (2020SciPy-NMeth).